Convert the following credit card rate to the APR. Nebraska, 0.02192% daily rate. (Round your answer to the nearest whole number.)
This one is difficult I got 21.9 is what I thought it was but I was wrong! Can you help me with this portion? Thanks!

Answer by stanbon(75887)

Convert the following credit card rate to the APR. Nebraska, 0.02192% daily rate. (Round your answer to the nearest whole number.)
-----------------------
APR/365 = 0.02192/100 = 0.0002192
---
APR = 365*0.0002192 = 0.08 = 8%
---------------
Cheers,
Stan H.